# Impact of Weak Lensing Mass Mapping Algorithms on Cosmology Inference

### v1

@@ A.2 Iterative Kaiser-Squires with DCT Inpainting (iKS)
where $p=0$ or $1$, $\Phi$ is the discrete cosine transform (DCT), and $\lambda$ is a Lagrangian parameter. The solution can be obtained using the forward-backward algorithm:
@@
where and ${\bm{\Delta}}_{{\Phi},\lambda}$ is the proximal operator defined in Pires et al. (2009), which consists of applying a DCT transform to $u$, thresholding the DCT coefficients, reconstructing an image from the thresholded coefficients, and normalizing the values of the filled-in pixels so that their standard deviation matches the standard deviation of the observed parts. This normalization can also be performed at every scale of the wavelet transform of the solution. Areas for which we possess information are processed as in the iterative KS case, while the inpainting regularisation affects areas with missing data (i.e. when $\textbf{M}=0$).
@@ A.3 MCALens
MCALens method was developed in Starck et al. (2021) to combine the advantages of two different methods with complementary strengths: the Wiener filter and sparsity-based methods. The Wiener filter models the convergence field as a Gaussian random field. It effectively reconstructs the large-scale Gaussian features of the convergence field, but it is sub-optimal in reconstructing the small-scale features, such as peaks. On the other hand, methods based on sparsity priors, e.g. Lanusse et al. (2016), assume that the convergence field is sparse when represented in a certain basis. Sparsity-based methods excel in capturing these peak structures, but may struggle to retrieve the Gaussian-like statistical properties that emerge on larger scales.
@@
MCALens then estimates $\bm{\kappa}_{G}$ and $\bm{\kappa}_{S}$ by applying the Morphological Component Analysis (MCA) algorithm, which is designed to separate different components mixed within a single signal or image, provided these components have distinct morphological properties. Specifically, MCA takes advantage of the sparsity of the different components in different dictionaries, which in the case of MCALens are a weighted Fourier basis for the Gaussian component,
@@
and a wavelet basis for the sparse component,
@@
The MCA algorithm is then solving the following optimization problem:
@@
where $\mathcal{C}_{G}$ and $\mathcal{C}_{S}$ are regularization terms for the Gaussian and sparse components, respectively. The solution to this problem is found by iteratively applying the following steps:
@@
    Estimate $\bm{\kappa}_{G}$ by assuming that $\bm{\kappa}_{S}$ is known, minimizing:
@@
    Estimate $\bm{\kappa}_{S}$ by assuming that $\bm{\kappa}_{G}$ is known, minimizing:
@@
To estimate the Gaussian component, the algorithm uses the Wiener filter, performing an iteration of the Proximal Wiener Filtering algorithm, as described in Starck et al. (2021), Section 2, at each MCA step. The Proximal Wiener Filtering is a variant of the standard Wiener filter (Bobin et al. 2012) which is designed to handle non-stationary noise, through the use of a Forward-Backward iterative algorithm. Thus, the penalty term $\mathcal{C}_{G}$ is in this case:
@@
For the sparse component, the algorithm cannot follow the typical $\ell_{1}$ of $\ell_{0}$-norm regularization used in the standard sparse recovery algorithms, as the large-scale wavelet coefficients overlap with low-frequency Fourier coefficients of the Gaussian field, complicating their separation. Instead, the algorithm adopts an alternative two-step approach:
@@
    Estimate the sparse component by solving the following minimization problem:
@@
This approach offers the advantage that when $\Omega$ is fixed, the algorithm behaves almost linearly, with the only nonlinearity being the positivity constraint imposed on $\bm{\kappa}_{S}$. This allows for the straightforward derivation of an approximate error map by simply propagating noise and relaxing the positivity constraint. The positivity constraint ensures that the reconstructed peaks are positive, which is not guaranteed in general, since peaks in the convergence field can be on top of voids and thus have negative values. The higher the non-Gaussianity of the convergence field, the more MCALens is expected to outperform linear methods like the Wiener filter.
@@ Appendix B Inference with simulated observations
In this appendix, we present the results of the inference analysis using as mock data the forward modelled HOS from the cosmo-SLICS simulations at the fiducial ($\Lambda$CDM) cosmology. The forward model applied to generate these fiducial data vectors is the same as the one used to generate the data vectors for the emulator training, described in Sections 4.1-4.3.
We then use the fiducial data vectors to estimate the size of the emulator’s interpolation error that was not captured by our covariance matrix. To measure this, we train an additional GPR model on the mean values of the HOS histograms across realizations at the 25 cosmologies. From this model, we obtain the prediction uncertainties from the GPR at the fiducial cosmology. We then incorporate these interpolation errors into our error budget by adding them to the diagonal of the covariance matrix of the HOS data vectors, following Heydenreich et al. (2021).
@@
The results for wavelet peak counts are shown in Fig. 11. These constraints were obtained using only 3 different scales: $16^{\prime},32^{\prime},\textrm{coarse}$. Including additional scales overly biased the GPR predictions due to the small number of cosmologies in the training set, leading to unreliable results. As in the case of single-scale peak counts, the resulting constraints are larger than those obtained using the emulator predictions as mock data, shown in Fig. 7 (both due to the use of only 3 scales, and the inclusion of the GPR interpolation error) but the improvement in the constraining power of MCALens over KS and iKS is still evident.
@@
Figure 11: Constraints on the parameters $\Omega_{\textrm{m}}$, $\sigma_{8}$, $h$, and $w_{0}$ from wavelet peak counts with using the $[16^{\prime},32^{\prime},\textrm{coarse}]$ scales, for simulated fiducial histograms. The contour lines are described in Fig. 6.

### v2

@@ A.2 Iterative Kaiser-Squires with DCT inpainting
where $p=0$ or $1$, $\bm{\Phi}$ is the DCT, and $\lambda$ is a Lagrangian parameter. The solution can be obtained using the forward-backward algorithm:
@@
where ${\bm{\Delta}}_{{\Phi},\lambda}$ is the proximal operator defined in Pires et al. (2009), which consists of applying a DCT transform to $\bm{t}$, thresholding the DCT coefficients, reconstructing an image from the thresholded coefficients, and normalising the values of the filled-in pixels so that their standard deviation matches the standard deviation of the observed parts. This normalisation can also be performed at every scale of the wavelet transform of the solution. Areas for which we possess information were processed as in the iterative KS case, while the inpainting regularisation affects areas with missing data (i.e. when $\textbf{M}=0$).
@@ A.3 MCALens
MCALens method was developed in Starck et al. (2021) to combine the advantages of two different methods with complementary strengths: the Wiener filter and sparsity-based methods. The Wiener filter models the convergence field as a Gaussian random field. It effectively reconstructs the large-scale Gaussian features of the convergence field, but it is suboptimal in reconstructing the small-scale features, such as peaks. On the other hand, methods based on sparsity priors (e.g. Lanusse et al. 2016) assume that the convergence field is sparse when represented in a certain basis. Sparsity-based methods excel in capturing these peak structures but can struggle to retrieve the Gaussian-like statistical properties that emerge on larger scales.
@@
MCALens then estimates $\bm{\kappa}_{G}$ and $\bm{\kappa}_{S}$ by applying the MCA algorithm, which is designed to separate different components mixed within a single signal or image, provided these components have distinct morphological properties. Specifically, MCA takes advantage of the sparsity of the different components in different dictionaries, which in the case of MCALens are a weighted Fourier basis $\mathbf{\Phi}_{G}$ for the Gaussian component,
@@
and a wavelet basis $\mathbf{\Phi}_{S}$ for the sparse component,
@@
The MCA algorithm is then solving the following optimisation problem:
@@
where $\mathcal{C}_{G}$ and $\mathcal{C}_{S}$ are regularisation terms for the Gaussian and sparse components, respectively. The solution to this problem is found by iteratively applying the following steps:
@@
    Estimate $\bm{\kappa}_{G}$ by assuming that $\bm{\kappa}_{S}$ is known, minimising:
@@
    Estimate $\bm{\kappa}_{S}$ by assuming that $\bm{\kappa}_{G}$ is known, minimising:
@@
To estimate the Gaussian component, the algorithm uses the Wiener filter, performing an iteration of the proximal Wiener filtering algorithm, as described in Starck et al. (2021), Section 2, at each MCA step. The proximal Wiener filtering is a variant of the standard Wiener filter (Bobin et al. 2012) which is designed to handle non-stationary noise, through the use of a Forward-Backward iterative algorithm. Thus, the penalty term $\mathcal{C}_{G}$ in this case is
@@
For the sparse component, the algorithm cannot follow the typical $\ell_{1}$ of $\ell_{0}$-norm regularisation used in the standard sparse recovery algorithms, as the large-scale wavelet coefficients overlap with low-frequency Fourier coefficients of the Gaussian field, complicating their separation. Instead, the algorithm adopts an alternative two-step approach:
@@
    Estimate the sparse component by solving the following minimisation problem:
@@
This approach offers the advantage that when $\Omega$ is fixed, the algorithm behaves almost linearly, with the only non-linearity being the positivity constraint imposed on $\bm{\kappa}_{S}$. This enables the straightforward derivation of an approximate error map by simply propagating noise and relaxing the positivity constraint. The positivity constraint ensures that the reconstructed peaks are positive, which is not guaranteed in general, since peaks in the convergence field can be on top of voids and thus have negative values. The higher the non-Gaussianity of the convergence field, the more MCALens is expected to outperform linear methods like the Wiener filter.
@@ Appendix B Inference with simulated observations
In this appendix, we present the results of the inference analysis using as mock data the forward-modelled HOS from the cosmo-SLICS simulations at the fiducial ($\Lambda$CDM) cosmology. The forward model applied to generate these fiducial data vectors is the same as the one used to generate the data vectors for the emulator training, described in Sections 4.1-4.3.
We then used the fiducial data vectors to estimate the size of the emulator’s interpolation error that was not captured by our covariance matrix. To measure this, we trained an additional GPR model on the mean values of the HOS histograms across realisations at the 25 cosmologies. From this model, we obtained the prediction uncertainties from the GPR at the fiducial cosmology. We then incorporated these interpolation errors into our error budget by adding them to the diagonal of the covariance matrix of the HOS data vectors, following Heydenreich et al. (2021).
@@
The results for wavelet peak counts are shown in Fig. 11. These constraints were obtained using only three different scales: $16^{\prime},32^{\prime},\textrm{coarse}$. Including additional scales overly biased the GPR predictions due to the small number of cosmologies in the training set, leading to unreliable results. As in the case of single-scale peak counts, the resulting constraints are larger than those obtained using the emulator predictions as mock data, shown in Fig. 7 (both due to the use of only three scales, and the inclusion of the GPR interpolation error) but the improvement in the constraining power of MCALens over KS and iKS is still evident.
@@
Figure 11: Constraints on the parameters $\Omega_{\textrm{m}}$, $\sigma_{8}$, $h$, and $w_{0}$ from wavelet peak counts using the $[16^{\prime},32^{\prime},\textrm{coarse}]$ scales, for simulated fiducial histograms. The contour lines are described in Fig. 6.